OpenStudy (miwqueen):

the range of the function f(x) = 7+3^x is a set of all positive numbers greater than. true or false

OpenStudy (miwqueen):

than 7

OpenStudy (mathmale):

Why not graph y=3^x? Then graph y=7+3^x. Doing this should help you answer this question. What does "range" mean in this context?

OpenStudy (sshayer):

\[if~x < 0,,3^x,Let~x=-y,y>0,3^x=3^{-y}=\frac{ 1 }{ 3^y }\rightarrow0~as y~ \rightarrow \infty\] hence \[f(x)\rightarrow7~as~x \rightarrow - \infty\] if \[x \geq 0,f(x)>7\] Range~is~[7,infinity)